OpenStudy (anonymous):

–3.62 – 4.7 A. –8.32 B. –1.08 C. 1.08 D. 8.32

OpenStudy (butterflydreamer):

Here's a helpful hint: -a - b = - (a + b) So, –3.62 – 4.7 = - ( 3.67 + 4.7) = ?

OpenStudy (anonymous):

just add the two numbers and then make it negative

OpenStudy (anonymous):

The correct answer is A. -8.32

OpenStudy (butterflydreamer):

@Stewpeas_0831 , please don't give out direct answers :)

OpenStudy (butterflydreamer):

@DIDITHEDAD ... two negatives don't make a positive unless you are multiplying. but we're subtracting. So we are just subtracting two numbers. SO to work it out: –3.62 – 4.7 = - ( 3.62 + 4.7) So let's work out what's inside the brackets. "( 3.62 + 4.7) " = 8.32 Now we know from earlier that –3.62 – 4.7 = - ( 3.62 + 4.7) = - (8.32) = -8.32

OpenStudy (anonymous):

oh, i knew that. sorry about that, i was thinking that it said 3.62 - (-4.7) sorry! @butterflydreamer is correct. :)

OpenStudy (butterflydreamer):

it's okaaay ^_^ !